# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 4.1. Generic smoothness
Example 27 (Singularity can happen at a special point in Part (i) of Theorem 3).
Let $V={\mathbb{C}}\langle x,y,z,w\rangle\supset W={\mathbb{C}}\langle x,y,z\rangle$ and $f=x^{2}y^{2}+z^{4}$ be a form of degree $4$. Then, $f$ represents a point in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$. Note that $\operatorname{rank}\phi_{2,2}(f)=4>3$ where $\phi_{a,d-a}:S^{d}V\to S^{a}V\otimes S^{d-a}V$ is the symmetric flattening. Theorem 3 (i) shows that a general form in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$ is a smooth point. But, here we show that $f$ is a singular point of $\sigma_{4}(v_{4}({\mathbb{P}}V))$. We know that the form $f_{D}=x^{2}y^{2}$ has Waring rank 3 so that $f_{D}=\ell_{1}^{4}+\ell_{2}^{4}+\ell_{3}^{4}$ for some $\ell_{i}\in{\mathbb{C}}[x,y]_{1}$. By [16, theorem 2.1] $f_{D}$ is also a singular point of $\sigma_{3}(v_{4}({\mathbb{P}}V))$. Since $f\in\langle f_{D},z^{4}\rangle$, by Terracini’s lemma, we see that $\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ and $\mathbb{T}_{\ell_{i}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ for any $i$. Further, because $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle v_{4}({\mathbb{P}}^{1})\rangle$ and $f_{D}$ has 1-dimensional secant fiber in its incidence, one can move $\ell_{i}$ along this ${\mathbb{P}}^{1}$. Thus, we have
(19) $\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\supseteq\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V),\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\rangle~{}.$
Note that using the parameterization (11) we can estimate the dimension of right hand side of (19). Take an affine open subset $\{[1:t:u_{1}:u_{2}]\}$ of ${\mathbb{P}}^{3}$ and (with a change of coordinates) let $z^{4}$ be $[1:0:1:1]$ and $\ell_{i}\in{\mathbb{P}}^{1}$ be reprsented by $[1:t:0:0]$ for $t\in{\mathbb{C}}$. Then, by (12) the embedded tangent space to $v_{4}({\mathbb{P}}V)$ at $[1:t:u_{1}:u_{2}]$ is given as the row span of
@@ 4.2. Singularity
In this subsection we will prove Part (ii) and (iii) both in Theorem 2 and Theorem 3, which show singularity of the $m$-subsecant loci $\Sigma_{k,d}(m)$ in the $k$-secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $\Sigma_{k,d}(m)$ is the union of all the $m$-subsecant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as (2), it is enough to prove the statements for $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
@@
As we noted above, here it is enough to show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for each $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. We will first prove $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ under the condition in Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$, next for Theorem 3 with $(k,d,m)\neq(8,4,3)$, and finally for the remaining $3$ cases $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$. Note that in the remaining $3$ cases, $(k,d,m)$ satisfy that $\beta=\binom{m+d}{m}-1\not<km+k-1$ and $k=\bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}-1\bigl{\rceil}$. Basically, we use the same idea for the proof, but a detailed way of estimation will be slightly different according to each case (due to defectivity and identifiability). The non-triviality of the singular locus, i.e. $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ can be directly obtained at the end by Lemma 10.
@@
$\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))~{}.$
@@
(20) $\langle\,\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
First of all, let us consider (ii) of Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$. Then $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ as in Remark 16. We have ${\mathbb{P}}^{\beta}=\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))$, since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills the whole ${\mathbb{P}}^{\beta}$. It is enough to discuss the following cases: (a1) $(k-1)m+k<\binom{m+d}{m}$; (a2) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$; (a3) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)=(3,2)$.
For the case (a1) (i.e., $(k-1)m+k<\binom{m+d}{m}$), by applying Proposition 15 (i) and Proposition 22 to $A=v_{d}^{-1}(q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\})$ in ${\mathbb{P}}^{m}$ and $\Lambda={\mathbb{P}}^{\beta}$, the number (14), which is a lower bound for the dimension of the left hand side in (20), is greater than or equal to $\beta+(n-m)(1+k+(km+k-1)-\beta)$. Thus, from the containment (20), we obtain
$\beta+(n-m)(1+k+(km+k-1)-\beta)\leq kn+k-1~{},$
@@
Now, assume $(k-1)m+k=\binom{m+d}{m}$ (equivalently, $km+k-1-\beta=m$). Then, in the same way as above, using Propositions 15 (ii) and taking $A={\mathbb{P}}^{m}=v_{d}^{-1}(Z)$ and $\Lambda=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta}$, we have
@@
For $(d,m)\neq(3,2)$ (i.e., the case of (a2)), by Propositions 15 (ii), it holds $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle\geq k+m$. Then $(n-m)(m+1)\leq km+k-1-\beta=m$, contrary to $n>m$. For $(d,m)=(3,2)$ (i.e., the case (a3)), we get $\beta=9$, $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle=5$, and $k=\lceil\frac{\binom{m+d}{m}}{m+1}\rceil=4$. The condition $(k,d,n)\neq(4,3,3)$ implies $n\geq 4$. Then we also have a contradiction since (21) does not hold. Hence we show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
Secondly, let us regard (ii) of Theorem 3. Except $(k,d,m)=(8,4,3)$, this part consists of (b1) $(k,d,m)=(7,3,4),(5,4,2),(9,4,3)$, and $(14,4,4)$ (i.e., the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ being defective); (b2) $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$ (i.e., just after defective case). By the same reason, we also have the containment (20) for these cases provided that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
For the defective cases (i.e., the case (b1)), it is known that all the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces in ${\mathbb{P}}^{\beta}$. So, as taking $A=v_{d}^{-1}(q_{i}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$ corresponding to the entry locus of $a$, by Proposition 22 a containment of the same kind as (20) implies
(22) $\beta-1+(n-m)(1+\dim\langle\,v_{d-1}(A)\,\rangle)\leq kn+k-1~{},$
@@
$n-m\leq\frac{\delta}{1+\dim\langle\,v_{d-1}(A)\,\rangle-k}~{}\bigg{(}\leq\frac{\delta}{1+\delta}<1\bigg{)}~{},$
@@
For the case (b2) i.e., just after defective cases $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$, we can also get a contradiction in a similar way. Since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))$ is a hypersurface in ${\mathbb{P}}^{\beta}$ in all these cases, by Lemma 14 we have $q_{i}(F)=v_{d}({\mathbb{P}}^{m})$ for any maximal irreducible component $F$ of $p^{-1}(a)$ so that we can take $A={\mathbb{P}}^{m}$. Further we see that $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is non-defective for any $n>m$ in each case and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ so that $\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))={\mathbb{P}}^{\beta}$. By Proposition 22, the containment (20) implies
$\beta+(n-m)(1+\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle)=\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}\leq kn+k-1~{},$
@@
which fails to hold in each 4 just after defective case above, because $n\geq m+1$. Thus, we obtain $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
For the remaining $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$, we first note that these are the only generic non-identifiability cases of subgeneric rank in which a generic secant fiber $p^{-1}(a)$ consists of not one, but two points (see [9, theorem 1.1]). Even though they do not satisfy ($\frac{{m+d\choose m}}{m+1}<k<\frac{{m+d\choose m}}{m+1}+1$ as above (so, not having a positive dimensional secant fibers), we can still get a proof by contradiction using another estimation.
@@
Similarly, suppose that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and take a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ so that $a$ is a smooth point both in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In each 3 cases of $(k,d,m)$, let $x_{1},\dots,x_{k}$ and $y_{1},\dots,y_{k}$ be the two secant fibers of $a$ (i.e., $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$ and $a\in\langle\,y_{1},\dots,y_{k}\,\rangle$) and let $x_{i}^{\prime}\in{\mathbb{P}}^{m}$ (resp. $y_{j}^{\prime}$) be the preimage of $x_{i}$, that is $v_{d}(x_{i}^{\prime})=x_{i}$, (resp. of $y_{j}$ with $v_{d}(y_{j}^{\prime})=y_{j}$). Now, let us set $A=\{x_{1}^{\prime},\dots,x_{k}^{\prime},y_{1}^{\prime},\dots,y_{k}^{\prime}\}$. Then, similarly we have a containment as before
@@
(23) $\big{\langle}\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\big{\rangle}\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
Since $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is general, we may assume that at least one $k$ points set of the secant fiber (say $x_{1}^{\prime},\dots,x_{k}^{\prime}$) are general points in ${\mathbb{P}}^{m}$. Then $\langle\,v_{d-1,m}(A)\,\rangle$ contains the $(k-1)$-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$. It follows that $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k$; otherwise, $\dim\langle\,v_{d-1,m}(A)\,\rangle\leq k-1$ implies $\langle\,v_{d-1,m}(A)\,\rangle=\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$ but by the generalized trisecant lemma ([28, proposition 1.4.3]) $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle\cap v_{d-1}({\mathbb{P}}^{m})=\{v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\}$ as a scheme, which gives a contradiction.
@@
$=kn+k+(n-m-1)~{},$
which is a contradiction since $n-m-1\geq 0$. Thus, it also holds that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ in these 3 generic non-identifiable cases.

### v3

@@ 4.1. Generic smoothness
Example 27 (Singularity can occur at a special point in Part (i) of Theorem 3).
Let
$V={\mathbb{C}}\langle x,y,z,w\rangle\supset W={\mathbb{C}}\langle x,y,z\rangle$
and $f=x^{2}y^{2}+z^{4}$ be a form of degree $4$. Then, $f$ represents a point in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$. Note that $\operatorname{rank}\phi_{2,2}(f)=4>3$ where $\phi_{a,d-a}:S^{d}V\to S^{a}V\otimes S^{d-a}V$ is the symmetric flattening. Theorem 3 (i) shows that a general form in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$ is a smooth point. But, here we show that $f$ is a singular point of $\sigma_{4}(v_{4}({\mathbb{P}}V))$. We know that the form $f_{D}=x^{2}y^{2}$ has Waring rank 3 so that $f_{D}=\ell_{1}^{4}+\ell_{2}^{4}+\ell_{3}^{4}$ for some $\ell_{i}\in{\mathbb{C}}[x,y]_{1}$. By [Han18, Theorem 2.1] $f_{D}$ is also a singular point of $\sigma_{3}(v_{4}({\mathbb{P}}V))$. Since $f\in\langle f_{D},z^{4}\rangle$, by Terracini’s lemma, we see that $\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ and $\mathbb{T}_{\ell_{i}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ for any $i$. Further, because $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle v_{4}({\mathbb{P}}^{1})\rangle$ and $f_{D}$ has 1-dimensional secant fiber in its incidence, one can move $\ell_{i}$ along this ${\mathbb{P}}^{1}$. Thus, we have
(27) $\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\supseteq\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V),\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\rangle.$
Note that using the parameterization (16) we can estimate the dimension of right hand side of (27). Take an affine open subset $\{[1:t:u_{1}:u_{2}]\}$ of ${\mathbb{P}}^{3}$ and (with a change of coordinates) let $z^{4}$ be $[1:0:1:1]$ and $\ell_{i}\in{\mathbb{P}}^{1}$ be represented by $[1:t:0:0]$ for $t\in{\mathbb{C}}$. Then, by (17) the embedded tangent space to $v_{4}({\mathbb{P}}V)$ at $[1:t:u_{1}:u_{2}]$ is given as the row span of
@@ 4.2. Singularity
In this subsection we will prove parts (ii) and (iii) both in Theorem 2 and Theorem 3, which show the singularity of the $m$-subsecant loci $\Sigma_{k,d}(m)$ in the $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $\Sigma_{k,d}(m)$ is the union of all the $m$-subsecant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as (3), it is enough to prove the statements for any $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
@@
As we noted above, it is enough here to show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for each $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
We will first prove that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ under the condition in Theorem 2 with $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$, next for Theorem 3 with $(k,d,m)\neq(9,3,5),(8,4,3),(9,6,2)$, and finally for $\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$ or $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$. Basically, we use the same idea for the proof, but a detailed way of estimation will be slightly different according to each case (due to secant defectivity and non-identifiability). The non-triviality of the singular locus, i.e., $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ can be directly obtained at the end by Lemma 10.
@@
$\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
@@
(28) $\bigl{\langle}\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\,\cup\bigcup_{x\in q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\bigr{\rangle}\,\subset\,\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
First of all, let us consider (ii) of Theorem 2 with $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$. Set $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$. Then $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ as in Remark 16. We have ${\mathbb{P}}^{\beta}=\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))$ since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills up the whole ${\mathbb{P}}^{\beta}$. It is enough to discuss the following three cases
  1.  (a1)
    $(k-1)m+k<\binom{m+d}{m}$,
  2.  (a2)
    $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$,
  3.  (a3)
    $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)=(3,2)$.
For the case (a1) (i.e., $(k-1)m+k<\binom{m+d}{m}$), we take $A=v_{d}^{-1}(q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\})$ in ${\mathbb{P}}^{m}$ and $\Lambda={\mathbb{P}}^{\beta}$. From Proposition 15 (i), it follows $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k+(km+k-1)-\beta$ for the ($d-1$)-uple Veronese embedding $v_{d-1,m}$ of ${\mathbb{P}}^{m}$. From Proposition 22, the dimension of the left hand side in (28) is greater than or equal to the number (19), which is
$\dim\langle\,\Lambda\cup v_{d}(A)\,\rangle+(n-m)\big{\{}1+\dim\langle\,v_{d-1,m}(A)\,\rangle\big{\}}\geq\beta+(n-m)(1+k+(km+k-1)-\beta).$
From the inclusion (28), we obtain
$\beta+(n-m)(1+k+(km+k-1)-\beta)\leq kn+k-1,$
@@
Now, assume $(k-1)m+k=\binom{m+d}{m}$ (equivalently, $km+k-1-\beta=m$). Then, in the same way as above, using Proposition 15 (ii) and taking $A={\mathbb{P}}^{m}=v_{d}^{-1}(Z)$ and $\Lambda=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta}$, we have
@@
For $(d,m)\neq(3,2)$ (i.e., the case of (a2)), Proposition 15 implies that $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle\geq k+m$. Then $(n-m)(m+1)\leq km+k-1-\beta=m$, contrary to $n>m$.
For $(d,m)=(3,2)$ (i.e., the case (a3)), we get $\beta=9$, $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle=5$, and $k=\lceil{\binom{m+d}{m}}/({m+1})\rceil=4$. The condition $(k,d,n)\neq(4,3,3)$ implies $n\geq 4$. Then we also have a contradiction since (29) does not hold. Hence we show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
Secondly, let us regard (ii) of Theorem 3. For $(k,d,m)=(10,3,5),(10,6,2)$, we have the same result as Theorem 2 since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ and it satisfies (a1),(a2) respectively. Then, except $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$, the remaining part of (ii) of Theorem 3 consists of the following two cases
  1.  (b1)
    $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$ (i.e., the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ being defective),
  2.  (b2)
    $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$ (i.e., just after the defective case).
By the same reason, we also have the inclusion (28) for these cases provided that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
For the case (b1), i.e., the defective case, it is known that all the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces in ${\mathbb{P}}^{\beta}$ (see [AH95]). So, as taking $A=v_{d}^{-1}(q_{i}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$ corresponding to the entry locus of $a$, by Proposition 22 a inclusion of the same kind as (28) implies
(30) $\beta-1+(n-m)(1+\dim\langle\,v_{d-1}(A)\,\rangle)\leq kn+k-1,$
@@
$n-m\leq\frac{\delta}{1+\dim\langle\,v_{d-1}(A)\,\rangle-k}\leq\frac{\delta}{1+\delta}<1,$
@@
For the case (b2), i.e., just after the defective case (b1), the $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills up ${\mathbb{P}}^{\beta}$, and hence $\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))={\mathbb{P}}^{\beta}$. Then, we can also get a contradiction in a similar way, as follows. Since the $(k-1)$-secant variety $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))$ is a hypersurface in ${\mathbb{P}}^{\beta}$, by Lemma 14 we have $q_{i}(F)=v_{d}({\mathbb{P}}^{m})$ for an irreducible component $F$ of $p^{-1}(a)$ for general $a\in{\mathbb{P}}^{\beta}$ so that we can take $A={\mathbb{P}}^{m}$. By Proposition 22, the inclusion (28) implies
$\beta+(n-m)(1+\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle)=\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}\leq kn+k-1,$
which fails to hold in (b2); more precisely, for $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$, the value $\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}-(kn+k-1)$ is equal to
$7n-33,4n-11,10n-35,20n-85,$
@@
respectively, which must be greater than 00 because of the condition $n\geq m+1$. Thus, we obtain $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
Now, we discuss the following two cases
  1.  (c1)
    $k=\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$ of Theorem 2; then, since we exclude $(k,d,m)=(5,3,3),(7,5,2)$, a general point $a\in{\mathbb{P}}^{\beta}=\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is not $k$-identifiable and the secant fiber $p^{-1}(a)$ consists of two or more points (see [GM17, Theorem 1]),
  2.  (c2)
    $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$ of Theorem 3; then, a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is not $k$-identifiable and $p^{-1}(a)$ consists of two points (see [COV17, Theorem 1.1]).
In these cases, even though they do not have positive dimensional secant fibers, we can still get a proof by contradiction using a different estimate, as follows.
@@
Similarly, suppose that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and take a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ so that $a$ is a smooth point in both $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. We take general $k$ points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ with $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$. By the non-identifiability, we have another set of $k$ points $y_{1},\dots,y_{k}\in v_{d}({\mathbb{P}}^{m})$ with $a\in\langle\,y_{1},\dots,y_{k}\,\rangle$ such that $(a,x_{1},\dots,x_{k})$ and $(a,y_{1},\dots,y_{k})$ are distinct in the secant fiber $p^{-1}(a)\subset I\subset{\mathbb{P}}^{\beta}\times(v_{d}({\mathbb{P}}^{m}))^{k}$ (modulo permutation on $(v_{d}({\mathbb{P}}^{m}))^{k}$). Let $x_{i}^{\prime}\in{\mathbb{P}}^{m}$ (resp. $y_{j}^{\prime}$) be the preimage of $x_{i}$, that is $v_{d}(x_{i}^{\prime})=x_{i}$ (resp. of $y_{j}$ with $v_{d}(y_{j}^{\prime})=y_{j}$).
Setting $A=\{x_{1}^{\prime},\dots,x_{k}^{\prime},y_{1}^{\prime},\dots,y_{k}^{\prime}\}$, we have an inclusion, similar to (28),
@@
(31) $\bigl{\langle}\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\,\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\bigr{\rangle}\,\subset\,\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
For the $(d-1)$-uple Veronese embedding $v_{d-1}=v_{d-1,m}:{\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta_{d-1}}$ with $\beta_{d-1}=\binom{m+d-1}{m}-1$, we have $\dim\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle=k-1$ since $x_{1}^{\prime},\dots,x_{k}^{\prime}$ are general in ${\mathbb{P}}^{m}$. The $(k-1)$-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$ is contained in $\langle\,v_{d-1,m}(A)\,\rangle$. On the other hand, the codimension of $v_{d-1}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta_{d-1}}$, that is $\binom{m+d-1}{m}-1-m$, is greater than or equal to $k$; it follows from (iii) of Lemma 17 in the case (c1), and from explicit calculations in the case (c2).
Then, we have $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k$, as follows. Otherwise, $\dim\langle\,v_{d-1,m}(A)\,\rangle\leq k-1$ implies $\langle\,v_{d-1,m}(A)\,\rangle=\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$. Since $y_{1}^{\prime},\dots,y_{k}^{\prime}\in A\subset{\mathbb{P}}^{m}$, it follows
$v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\in\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle\cap v_{d-1}({\mathbb{P}}^{m}),$
where the right hand side must be $\{v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\}$ because of the generalized trisecant lemma ([Rus16, Proposition 1.4.3]), which gives a contradiction.
@@
$=kn+k+(n-m-1),$
which is a contradiction since $n-m-1\geq 0$. Thus, it also holds that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ in these generic non-identifiable cases.
Note that for $(k,d,m)=(10,3,5),(9,4,3),(10,6,2)$, i.e., just after the non-identifiable case (c2), the singularity is already shown in the second part of this proof, where $(k,d,m)=(9,4,3)$ is also in the defective case (b1). The case $(k,d,m)=(5,3,3),(7,5,2)$, which is excluded from (c1), belongs to (i) of Theorem 3; in this sense, the non-trivial singularity does not appear for $(d,m)=(3,3),(5,2)$.